### v1

@@ Abstract
Measurement of lithium abundances in solar-type stars have shown that standard models of stellar evolution are incapable of explaining the observed depletion as a function of stellar age. Beryllium is one of the lightest elements that can be measured in stellar photospheres, and it can be burned in relatively low temperatures. Studying its abundances as a function of stellar age can provide important constraints to stellar mixing models, as the level of depletion as a function of time will indicate how deep the photospheric material must be dredged to explain the observed abundances of both elements. In an effort to provide the most stringent constraints for non-standard stellar mixing models, we observed a sample of solar-twins and concomitantly analyzed their lithium and beryllium abundances. Unlike what is typically observed for lithium, we find that beryllium does not decrease as a function of stellar age, constraining models that predict burning of both materials. Based on our data, models that invoke convective overshoot and convective settling are preferred over typical rotationaly-induced mixing models, as the later burn Be in excess while the former do not. Previous works also proposed mixing due to gravity waves as a possible explanation for observed abundances, which can fit our data as well. We also confirm previous finds of an increase in Be abundance as a function of metallicity, indicative of galactic production via cosmic ray spallation.
@@ 1 Introduction
The two lightest metals in the Universe are beryllium (Be) and lithium (Li). These fragile atoms are burned at relatively low temperatures ($2.5\times 10^{6}$ and $3.5\times 10^{6}$ K, Li and Be respectively). In the Sun, Li burning requires temperatures that are somewhat higher than those reached at the base of the convective zone, according to standard models of stellar evolution (e.g., Schwarzschild et al., 1957; Weymann & Sears, 1965; Pinsonneault, 1997). Nevertheless, the observed photospheric abundances of Li (estimated using the most up-to-date methodology, i.e., three dimensional and non-local thermodynamic equilibrium, non-LTE, modelling) are considerably lower than those inferred from the analysis of meteorites ($0.96\times 3.25$ for the Solar Photosphere vs. Meteorites, Greenstein & Richardson, 1951; Weymann & Sears, 1965; Grevesse, 2019; Asplund et al., 2021), which are representative of the Li composition of the Sun at the Zero Age Main Sequence (ZAMS). Therefore, there must have been lithium burning after the Sun entered the main-sequence. Which leads to the conclusion that there must be physical mechanisms responsible for transporting Li deeper than the base of the convective zone, where temperatures are high enough for it to be destroyed during the main-sequence.
Different mechanisms have been proposed to explain this “extra mixing”, such as rotation (e.g. Pinsonneault et al., 1989); internal gravity waves (e.g., Charbonnel & Talon, 2005); microscopic diffusion and gravitational settling (e.g., Michaud et al., 2004); convective overshooting (e.g., Xiong & Deng, 2007); and convective settling (e.g., Andrássy & Spruit, 2015). The mechanism(s) responsible for the observed extra mixing could be one of or, more likely, a combination of, these proposed candidates. This extra mixing is clearly observed in the analysis of solar twins, in which there is a correlation between lithium abundance and main-sequence lifetime (e.g., Carlos et al., 2016, 2019; Boesgaard et al., 2022; Martos et al., 2023). Lithium depletion have been observed in early main-sequence stars of other spectral types as well (e.g., Boesgaard & Tripicco, 1986). Thévenin et al. (2017) proposed that Li depletion in solar twins can be explained through pre-main sequence burning. However, Carlos et al. (2019), Martos et al. (2023) and Rathsam et al. (2023), using large samples of solar twins and solar analogs, confirmed the existence of a correlation between Li abundances and stellar ages for solar twins, a strong evidence of main-sequence depletion.
@@
Using beryllium as a secondary probe of stellar mixing is not new in the literature, an ongoing effort since the discovery by Boesgaard & Tripicco (1986) that star 110 Her is depleted in both Li and Be. We focus on two recent efforts: Tucci Maia et al. (2015) and Boesgaard et al. (2022). The first analyzed a small sample of solar-twins, and assumed a solar value of A(Be)=$1.38$ dex (as per Asplund et al., 2009). They found the Sun to be depleted in Be compared to their sample of 8 stars and found no signal of beryllium depletion as a function of age. However, the beryllium abundances they found seem unusually large. The second, and more recent effort, is a study analyzing beryllium abundances in a sample of approximately 50 one-solar-mass stars. They conclude that rotationally-induced extra mixing (e.g., Deliyannis & Pinsonneault, 1997; Pinsonneault et al., 2014), possibly the most popular among the explanations available in the literature to explain lithium depletion during the main-sequence, is not consistent with their data, as Be is not observed to be depleted. Moreover, Boesgaard et al. (2022) note that mixing due to gravity waves could be a better explanation, as it does not deplete Be as much as Li. They also find a correlation between beryllium and metallicity, a possible indication of galactic nucleosynthesis.
@@ 2.1 Target Selection and Observations
All our targets were observed with the Ultraviolet and Visual Echelle Spectrograph (UVES, Dekker et al., 2000), at the VLT telescope under program ID 111.24VX.001. We used the blue configuration with the CD1 dichroic, centered at $346$ nm, and the 0.7” slit, yielding a resolving power of R55,000absent55000$\approx$ 55,000≈ 55 , 000, with an effective wavelength range from $302.5$ to $388.4$ nm. A log of our observations showing the full target list, along with the spectral signal-to-noise ratio at $346$ nm is shown in Table 1. We also include in our analysis a solar spectrum, observed via the reflected light of the Juno asteroid (kindly provided by Jorge Meléndez, via private communication), using the same configuration in the blue, with the same resolving power and high S/N.
@@ 2.2 Data Reduction
We retrieved the processed data products from the ESO archive. These include bias subtraction, flat field correction, optimized order extraction, S/N calculation, and barycentric velocity calculation. We then used the spectrum analysis tool iSpec [^2] https://www.blancocuaresma.com/s/iSpec (Blanco-Cuaresma et al., 2014a; Blanco-Cuaresma, 2019a) to perform radial velocity corrections and spectra normalization. For radial velocity corrections we performed a cross-correlation analysis using the VALD [^3] http://vald.astro.uu.se/ linelist, which is integrated into iSpec, to find a suitable radial-velocity (RV) correction for our abundance analysis. We then reduced the wavelength range of our RV corrected spectra to be between $306$ and $318$ nm. The only observable beryllium lines are the $3130.42$ and $3131.06$ transitions.
@@
The spectral region with the two beryllium atomic transition contains a multitude of different atomic and molecular transitions. Therefore, the normalization process can be prone to errors. To mitigate uncertainties in our measurements due to the normalization process, we normalize only our wavelength region containing the beryllium lines, and we do this using the previously normalized Juno spectrum in the same region as a template. In Figure 1 we show the Juno normalized spectrum (black solid lines) overplot with the normalized spectrum of three stars with varying metallicities (HIP 64713, HIP 54287, and HIP 109821) to show that the continuum does not vary significantly in our sample. The process of normalizing all our science targets using an observed solar spectrum as a template guarantees that our differential analysis will contain the same (if any) systematics due to the normalization process. Therefore, we mitigate our relative uncertainties and guarantee that the relative abundances are precise.
To determine stellar parameters and lithium abundance, we employed stellar spectra collected by the HARPS (Mayor et al., 2003) spectrograph. This instrument is mounted on the 3.6 m telescope at La Silla observatory (Atacama/Chile) and delivers a resolving power of $R=115,000$ over a $\sim 380-690$ nm wavelength range. We retrieved reduced HARPS spectra from the ESO Archive [^4] http://archive.eso.org/wdb/wdb/adp/phase3_spectral/form?phase3_collection=HARPS. For each one of our targets, we selected the twentieth highest SNR spectra under the following observational constraint: lowest possibly air-mass ($\lesssim$ 2), and highest possible Moon angular separation ($\gtrsim$ 20 degrees) and SNR at 5500 Å($\gtrsim$ 100 pixels-1). In those cases where eventually at least one of such conditions could not be met, we gradually relaxed each one of the observational constraints ensuring a high quality master spectrum for each star.
@@ 3 Data Analysis
We determined atmospheric parameters through strictly differential analysis following (Martioli et al., 2023) (see their section 3.4 and Appendix B) and Lorenzo-Oliveira et al. (in prep). Here, we provide a brief description of the process. Equivalent widths (EWs) of the atomic transitions of Fe I, and Fe II(Meléndez et al., 2014) were measured using an automatic Python script that fits Gaussian to a given line profile bracketed by a 6 Å spectral window. The method is based on line-by-line equivalent width measurements between the Sun and the targets, choosing similar pseudo-continuum regions for both stars.
@@
We estimated ages, masses, and radii of our sample through a Bayesian approach anchored on a different set of structural models, Yonsei-Yale evolutionary tracks (Yi et al., 2001; Demarque et al., 2004). With the input parameters and its respective errors ($T_{\rm eff}$, $\rm{[Fe/H]}$, [$\alpha$/Fe], $\log g$, Gaia parallaxes, and $G$, $G_{\rm{BP}}$, and $G_{\rm{RP}}$ magnitudes), and posterior distribution functions derived from a proper likelihood marginalization along all possible evolutionary steps. Then, the likelihood is weighted by its respective mass (the Salpeter 1955 mass function) and metallicity (Casagrande, 2018) priors to obtain the desired evolutionary parameter’s median (50th percentile) and the ±$\pm 1\sigma$ intervals (16th to 84th percentile). For more details we refer to Lorenzo-Oliveira et al. (2019).
@@ 3.1 Lithium Abundance calculations
To determine $A$(Li) in our sample of solar twins, we followed the prescription given in Yana Galarza et al. (2016), but for HARPS spectra. In summary, we determine macroturbulence velocity using the relation found by dos Santos et al. (2016) and projected rotational velocities by fitting the line profiles of four iron lines (6027.050, 6151.618, 6165.360, 6705.102 Å,) and one nickel line (6767.772 Å). Then, we synthesized the region around the asymmetric 6707.75 Å Li line using the radiative transfer code MOOG (Sneden, 1973a), which assumes local thermodynamical equilibrium (LTE), with Kurucz model atmospheres (Castelli & Kurucz, 2003). We corrected all our non-LTE abundances using the grid from Lind et al. (2009).
@@ 3.2 Beryllium Abundance calculations
As already pointed out, the near ultraviolet (just ultraviolet from now-on) region containing the beryllium transitions is filled with different atomic and molecular transitions, which makes not only the normalization process difficult, but also complicates the process of generating the best match between the synthetic spectrum with the observed data. Further adding to this difficulty is the fact that not all the transitions in the region have atomic and molecular data measured in the laboratory, adding their uncertainties to the final systematic uncertainty of our data analysis. The beryllium transitions are also prone to departures from thermodynamical equilibrium (non-LTE or NLTE departures) which can be significant (e.g., Korotin & Kučinskas, 2022).
As it is out of the scope of our study to produce a full NLTE analysis of every spectrum, which would be time and computationally expensive, we decided to perform a differential approach to our analysis. We start by reproducing the observed solar spectrum with the current state-of-the-art solar abundances. We assume for all elements, except beryllium, the solar abundances from Asplund et al. (2021). For Be we assume the estimated abundance of Korotin & Kučinskas (2022). They showed in their work that their new calculation of the solar beryllium, using a NLTE methodology, yielded a solar photospheric beryllium abundance of A(Be)$=1.32\pm 0.05$ dex, fully compatible with the meteoritic values (A(Be)$=1.31\pm 0.04$ dex).
We adopt a solar macroturbulent velocity of $\rm{V_{mac}}=3.5\ \rm{km.s^{-1}}$, a rotational velocity of $\rm{vsin}i=1.9\ \rm{km.s^{-1}}$ (Bruning, 1984; Saar & Osten, 1997), a limb-darkening coefficient of $0.6$, and we calculated the instrumental broadening (assuming a Gaussian full-width half maximum - FWHM) at $3131$, for a resolving power of R$=55,000$, to be $0.06$.
@@
Our initial analysis did not provide a sufficiently good match between observed and synthesized spectrum, which we mostly attribute to two aspects: 1) The adopted solar abundances were mostly estimated using NLTE and, at times, 3D model atmospheres, opposed to our 1D, LTE analysis; 2) The shortcomings of analyzing such short wavelengths, namely normalization difficulties, and/or atomic/molecular line data of insufficient quality available.
@@
This updated linelist is used throughout our analysis. We warn the reader that, although we make our linelist available, it should be used with care for stars that are not like the Sun. It was produced for such a solar-twins analysis. By adapting the linelist in this way, we also guarantee that the analysis of our full sample will be done differentially, meaning that observed abundances will be estimated in relation to the Sun (as a typical differential analysis using equivalent widths would do). In practice, this will decrease our systematic uncertainties and increase our analysis precision (e.g., Reggiani & Meléndez, 2017; McKenzie et al., 2022).
We then synthesize the Be region of our entire sample using this updated linelist described above, using 1D MARCS model atmospheres and the LTE radiative transfer code MOOG. In Table 2 we present the Be abundances of our sample.

### v2

@@ Abstract
Measurements of lithium abundances in solar-type stars have shown that standard models of stellar evolution are incapable of explaining the observed depletion as a function of stellar age. Beryllium is one of the lightest elements that can be measured in stellar photospheres, and it can be burned in relatively low temperatures. Studying its abundances as a function of stellar age can provide important constraints to stellar mixing models, as the level of depletion as a function of time will indicate how deep the photospheric material must be dredged to explain the observed abundances. In an effort to provide the most stringent constraints for non-standard stellar mixing models, we observed a sample of solar-twins and concomitantly analyzed their lithium and beryllium abundances. Unlike what is typically observed for lithium, we found that beryllium does not decrease as a function of stellar age along the main-sequence, constraining models that predict burning of both materials. Based on our data, models that invoke convective overshoot and convective settling are preferred over typical rotationaly-induced mixing models, as the later burn Be in excess while the former do not. Previous works also proposed mixing due to gravity waves as a possible explanation for observed abundances, which can fit our data as well. Furthermore, based on our solar twins, Be depletion likely happens within the first $\sim 1$ Gyr. We also confirm previous findings of an increase in Be abundance as a function of metallicity, indicative of galactic production via cosmic ray spallation.
@@ 1 Introduction
The two lightest metals in the Universe are beryllium (Be) and lithium (Li). These fragile atoms are burned at relatively low temperatures ($2.5\times 10^{6}$ and $3.5\times 10^{6}$ K, Li and Be respectively). In the Sun, Li burning requires temperatures that are somewhat higher than those reached at the base of the convective zone, according to standard models of stellar evolution (e.g., Schwarzschild et al., 1957; Weymann & Sears, 1965; Pinsonneault, 1997). Nevertheless, the observed photospheric abundances of Li (estimated using the most up-to-date methodology, i.e., three-dimensional and non-local thermodynamic equilibrium, non-LTE, modeling) are considerably lower than those inferred from the analysis of meteorites ($0.96\times 3.25$ for the Solar Photosphere vs. Meteorites, Greenstein & Richardson, 1951; Weymann & Sears, 1965; Grevesse, 2019; Asplund et al., 2021), which are representative of the Li composition of the Sun at the Zero Age Main Sequence (ZAMS). Therefore, there must have been lithium burning after the Sun entered the main-sequence. Which leads to the conclusion that there must be physical mechanisms responsible for transporting Li deeper than the base of the convective zone, where temperatures are high enough for it to be destroyed during the main-sequence.
Different mechanisms have been proposed to explain this “extra mixing”, such as rotation (e.g. Pinsonneault et al., 1989); internal gravity waves (e.g., Charbonnel & Talon, 2005); microscopic diffusion and gravitational settling (e.g., Michaud et al., 2004); convective overshooting (e.g., Xiong & Deng, 2007; Zhang et al., 2019); and convective settling (e.g., Andrássy & Spruit, 2015). The mechanism(s) responsible for the observed extra mixing could be one of or, more likely, a combination of, these proposed candidates. This extra mixing is clearly observed in the analysis of solar twins, in which there is a correlation between lithium abundance and main-sequence lifetime (e.g., Carlos et al., 2016, 2019; Boesgaard et al., 2022; Martos et al., 2023). Lithium depletion has been observed in early main-sequence stars of other spectral types as well (e.g., Boesgaard & Tripicco, 1986). Thévenin et al. (2017) proposed that Li depletion in solar twins can be explained through pre-main sequence burning. However, Carlos et al. (2019), Martos et al. (2023) and Rathsam et al. (2023), using large samples of solar twins and solar analogs, confirmed the existence of a correlation between Li abundances and stellar ages ($\tau$) for solar twins, a strong evidence of main-sequence depletion.
@@
Using beryllium as a secondary probe of stellar mixing is not new in the literature, an ongoing effort since the discovery by Boesgaard & Tripicco (1986) that star 110 Her is depleted in both Li and Be. We focus on two recent efforts: Tucci Maia et al. (2015) and Boesgaard et al. (2022). The first analyzed a small sample of solar-twins, and assumed a solar value of A(Be)=$1.38$ dex (as per Asplund et al., 2009). They found the Sun to be depleted in Be compared to their sample of 8 stars and found no signal of beryllium depletion as a function of age. However, the beryllium abundances they found seem unusually large. The second, and more recent effort, is a study analyzing beryllium abundances in a sample of approximately 50 one-solar-mass stars. They concluded that one of the most prominent explanations in the literature to explain the lithium abundance depletion, rotationally-induced extra mixing (e.g., Deliyannis & Pinsonneault, 1997; Pinsonneault et al., 2014), is not consistent with their data, as Be is not sufficiently depleted. Boesgaard et al. (2022) claim that mixing due to gravity waves could be a potential explanation, as it does not deplete Be as much as Li. They also find a correlation between beryllium and metallicity, a possible indication of galactic nucleosynthesis.
@@ 2.1 Target Selection and Observations
First, we selected a sample of solar-twins by applying the photometric cuts proposed by Yana Galarza et al. (2021) to the Gaia database. We then searched archival high-resolution, high-signal-to-noise (S/N) spectroscopic data. Using the cross-match between photometrically-selected solar-twins and spectra availability, we created of bona-fide solar-twins with high precision photospheric stellar parameters. All of our targets have effective temperatures, surface gravities, metallicities, microturbulent velocities, as well as fundamental stellar parameters, namely masses and stellar ages. A description of our methodologies and the full sample is in preparation (Lorenzo-Oliveira et al. in prep).
@@
All our targets were observed with the Ultraviolet and Visual Echelle Spectrograph (UVES, Dekker et al., 2000), at the VLT telescope under program ID 111.24VX.001. We used the blue configuration with the CD1 cross-disperser, centered at $346$ nm, and the 0.7” slit, yielding a resolving power of R55,000absent55000$\approx$ 55,000≈ 55 , 000, with an effective wavelength range from $302.5$ to $388.4$ nm. A log of our observations showing the full target list, along with the spectral signal-to-noise at $346$ nm is shown in Table 1. We also include in our analysis a solar spectrum, observed via the reflected light of the Juno asteroid (kindly provided by Jorge Meléndez, via private communication), using the same configuration in the blue, with the same resolving power and high S/N.
@@ 2.2 Data Reduction
We retrieved the UVES processed data products from the ESO archive. These include bias subtraction, flat field correction, optimized order extraction, S/N calculation, and barycentric velocity calculation. We then used the spectrum analysis tool iSpec [^2] https://www.blancocuaresma.com/s/iSpec (Blanco-Cuaresma et al., 2014a; Blanco-Cuaresma, 2019a) to perform radial velocity corrections and spectra normalization. For radial velocity corrections we performed a cross-correlation analysis using the VALD [^3] http://vald.astro.uu.se/ linelist, which is integrated into iSpec, to find a suitable radial-velocity (RV) correction for our abundance analysis. We then reduced the wavelength range of our RV corrected spectra to be between $306$ and $318$ nm. The only observable beryllium lines are the $3130.42$ and $3131.06$.
@@
The spectral region containing the two beryllium atomic transition contains a multitude of different atomic and molecular transitions. Therefore, the normalization process can be prone to errors. To mitigate uncertainties in our measurements due to the normalization process, we normalize only our wavelength region containing the beryllium lines, and we do this using the previously normalized Juno spectrum in the same region as a template. In Figure 1 we show the Juno normalized spectrum (black solid lines) overplot with the normalized spectrum of three stars with varying metallicities (HIP 64713, HIP 54287, and HIP 109821) to show that the continuum does not vary significantly in our sample. The process of normalizing all our science targets using an observed solar spectrum as a template guarantees that our differential analysis will contain the same (if any) systematics due to the normalization process. Therefore, we mitigate our relative uncertainties and guarantee that the relative abundances are precise.
To determine stellar parameters and lithium abundance, we employed stellar spectra collected by the HARPS (Mayor et al., 2003) spectrograph. This instrument is mounted on the 3.6 m telescope at La Silla observatory (Atacama/Chile) and delivers a resolving power of $R=115,000$ over a $\sim 380-690$ nm wavelength range. We retrieved reduced HARPS spectra from the ESO Archive [^4] http://archive.eso.org/wdb/wdb/adp/phase3_spectral/form?phase3_collection=HARPS. For each one of our targets, we selected the twentieth highest SNR spectra under the following observational constraint: lowest possible air-mass ($\lesssim$ 2), and highest possible Moon angular separation ($\gtrsim$ 20 degrees) and SNR at 5500 Å($\gtrsim$ 100 pixels-1). In those cases where eventually at least one of such conditions could not be met, we gradually relaxed each one of the observational constraints, ensuring a high quality master spectrum for each star.
@@ 3 Data Analysis
3.1 Stellar Parameters
We determined atmospheric parameters through strictly differential analysis following Martioli et al. (2023) (see their section 3.4 and Appendix B) and Lorenzo-Oliveira et al. (in prep). Here, we provide a brief description of the process. Equivalent widths (EWs) of the atomic transitions of Fe I, and Fe II(Meléndez et al., 2014) were measured using an automatic Python script that fits Gaussian to a given line profile bracketed by a 6 Å spectral window. The method is based on line-by-line equivalent width measurements between the Sun and the targets, choosing similar pseudo-continuum regions for both stars.
@@ 3.1 Stellar Parameters
We estimated ages (here-forth $\tau_{\mathrm{iso}}$ for our inferred ages - based on isochrone fitting), masses, and radii of our sample through a Bayesian approach anchored on a different set of structural models, Yonsei-Yale evolutionary tracks (Yi et al., 2001; Demarque et al., 2004). With the input parameters and its respective errors ($T_{\rm eff}$, $\rm{[Fe/H]}$, [$\alpha$/Fe], $\log g$, Gaia parallaxes, and $G$, $G_{\rm{BP}}$, and $G_{\rm{RP}}$ magnitudes), and posterior distribution functions derived from a proper likelihood marginalization along all possible evolutionary steps. Then, the likelihood is weighted by its respective mass (the Salpeter 1955 mass function) and metallicity (Casagrande, 2018) priors to obtain the desired evolutionary parameter’s median (50th percentile) and the ±$\pm 1\sigma$ intervals (16th to 84th percentile). For more details we refer to Lorenzo-Oliveira et al. (2019).
@@ 3.2 Lithium Abundance calculations
To determine $A$(Li) in our sample of solar twins, we followed the prescription given in Yana Galarza et al. (2016), but for HARPS spectra. In summary, we determine macroturbulence velocity using the relation found by dos Santos et al. (2016) and projected rotational velocities by fitting the line profiles of five iron lines (6027.050, 6151.618, 6165.360, 6705.102 Å,) and one nickel line (6767.772 Å). Then, we synthesized the region around the asymmetric 6707.75 Å Li line using the radiative transfer code MOOG (Sneden, 1973a), which assumes local thermodynamical equilibrium (LTE), with Kurucz model atmospheres (Castelli & Kurucz, 2003). We corrected all our non-LTE (NLTE) abundances using the grid from Lind et al. (2009).
@@ 3.3 Beryllium Abundance calculations
As already pointed out, the near ultraviolet (just ultraviolet from now on) region containing the beryllium transitions is filled with different atomic and molecular transitions, which makes not only the normalization process difficult, but also complicates the process of generating the best match between the synthetic spectrum and the observed data. Further adding to this difficulty is the fact that not all the transitions in the region have atomic and molecular data measured in the laboratory, adding their uncertainties to the final systematic uncertainty of our data analysis. The beryllium transitions are also prone to departures from local thermodynamical equilibrium (non-LTE or NLTE departures) which can be significant (e.g., Korotin & Kučinskas, 2022; Amarsi et al., 2024).
As it is out of the scope of our study to produce a fully NLTE analysis of every spectrum, which would be time and computationally expensive, we decided to perform a differential approach to our analysis. We start by reproducing the observed solar spectrum with the current state-of-the-art solar abundances. We assume for all elements, except beryllium, the solar abundances from Asplund et al. (2021). For Be we performed two calibrations. First, we assumed the estimated abundance of Korotin & Kučinskas (2022). In our second iteration, we assumed the abundance from Amarsi et al. (2024). Korotin & Kučinskas (2022) used a 1D NLTE methodology estimated A(Be)$=1.32\pm 0.05$ dex for the present-day solar beryllium abundance, fully compatible with the meteoritic values of A(Be)$=1.31\pm 0.04$ dex. Based on 3D NLTE models, Amarsi et al. (2024) found slightly a lower abundance: A(Be)$=1.21\pm 0.05$ dex.
We adopt a solar macroturbulent velocity of $\rm{V_{mac}}=3.5\ \rm{km.s^{-1}}$, a rotational velocity of $\rm{vsin}i=1.9\ \rm{km.s^{-1}}$ (Bruning, 1984; Saar & Osten, 1997), a limb-darkening coefficient of $0.6$, and we calculate the instrumental broadening (assuming a Gaussian full-width half maximum - FWHM) at $3131$, for a resolving power of R$=55,000$, to be $0.06$.
@@
Our initial analysis did not provide a sufficiently good match between observed and synthesized spectrum, which we mostly attribute to two aspects: 1) The adopted solar abundances were mostly estimated using NLTE and, at times, 3D model atmospheres, opposed to our 1D, LTE analysis; 2) The shortcomings of analyzing such short wavelengths, namely normalization difficulties, and/or atomic/molecular line data of insufficient quality as well as the known issue of the missing oppacities in the Be region (e.g., Amarsi et al., 2024).
@@
We proceed our analysis with the linelist that reproduces the present-day solar beryllium as calculated by Amarsi et al. (2024) (A(Be)$=1.21$). Then, we used our linelist to synthesize the beryllium region for the stars in our sample. Because we are first fine-tuning the linelist to match the solar properties, the resulting abundance zero-point shared by the solar twins is A(Be)$=1.21$. In other words, we adopt a differential approach in respect to the Sun to derive Be abundances for our sample stars. This approach takes advantage of the known solar spectroscopic similarity with solar twins to improve measurements of minute abundance differences compared to the Sun’s absolute value. Regardless of the accuracy of the zero-point (A(Be)$=1.21$ or A(Be)$=1.32$), the precision on the relative abundances will be used throughout the analysis to study the depletion of Be as a function of time.
To synthesize the Be region in our sample of solar-twins we assumed a limb-darkening coefficient of $0.6$, and we calculate the instrumental broadening (assuming a Gaussian full-width half maximum - FWHM) at $3131$, for a resolving power of R$=55,000$, to be $0.06$. We calculated the macroturbulent velocity of each star following dos Santos et al. (2016). Rotational velocity (vsin(i)) was estimated along with the Be abundance at the region. We employed the 1D MARCS model atmospheres coupled to the LTE radiative transfer code MOOG, and the solar fine-tuned linelist described above. In Table 2 we present the Be abundances of our sample.